# Reducing Biases in Record Matching Through Scores Calibration

### v2

@@ 5       $\text{append}(\text{scores}_{g},s(p_{i})+\mathcal{N}(0,\sigma^{2}))$;
7sort($\text{sort}(\text{scores}_{a})$; $\text{sort}(\text{scores}_{b})$;
@@ 11       $\text{pos}_{g}\leftarrow\text{pos}_{g}+1$;
14 $\text{pos}_{\bar{g}}\leftarrow\left\lceil q\times|\text{scores}_{\bar{g}}|\right\rceil$; /* Cross-group map */
@@
The second point concerns the input dataset $D=\{p_{1},\dots,p_{|D|}\}$, used to estimate $P_{a}$ and $P_{b}$. This dataset does not require ground-truth labels–only matching scores are needed, which can be obtained by applying the score function $s$. As a result, $s$ can be calibrated as a black box without access to the training data used to learn $s$, such as in deep learning-based matchers. This significantly broadens the applicability of the calibration method. The only assumption is that the pairs in $D$ are i.i.d. samples from the space of possible record pairs and that the dataset is large enough to reliably estimate the score distributions. The primary goal of the algorithm is to calibrate the scores for one or more query pairs. When the number of query pairs is large enough (as discussed in Section IV-B), the dataset $D$ can simply be the set of query points itself. We examine the effect of choosing a dataset $D$ with a distribution different from that of the query points in Section VI-B5. We now describe each step of Algorithm 1.
@@
To calibrate the score for a query pair $p$, the algorithm identifies the group $g$ of $p$ (Line 1) and initializes a $pos$ variable within the sorted score list (Line 1). It then iterates through the sorted scores to locate the position where $s(p)$ would fit, counting scores greater than $s(p)$ within group $g$ (Lines 11). This position, relative to the list length, defines a quantile $q$, representing the percentage of scores in group $g$ greater than $s(p)$ (Line 1). The algorithm then calculates $\text{pos}_{\bar{g}}=\left\lceil q\times|\text{scores}_{\bar{g}}|\right\rceil$, where $\bar{g}$ is the opposite group (i.e., if $g=a$, then $\bar{g}=b$, and vice versa). This position is used to find the score in the other group’s list that corresponds to the same quantile $q$ as the query score in its own group.
@@
Consider dataset $D$ shown in Figure 2, containing 15 record pairs scored by $s$ and labeled a or b by a sensitive attribute. The pairs and scores are: $(0.45,a)$, $(0.82,a)$, $(0.90,b)$, $(0.71,a)$, $(0.84,b)$, $(0.67,a)$, $(0.38,b)$, $(0.98,b)$, $(0.36,b)$, $(0.38,a)$, $(0.32,b)$, $(0.29,a)$, $(0.24,b)$, $(0.21,b)$, $(0.19,b)$. Following the calibration algorithm, Gaussian noise $\mathcal{N}(0,0.05^{2})$ is added to each score, resulting in the following (rounded) values: $(0.46,a)$, $(0.80,a)$, $(0.89,b)$, $(0.72,a)$, $(0.85,b)$, $(0.65,a)$, $(0.37,b)$, $(0.97,b)$, $(0.35,b)$, $(0.39,a)$, $(0.31,b)$, $(0.28,a)$, $(0.25,b)$, $(0.22,b)$, $(0.18,b)$. Sorting these in descending order gives $\text{scores}_{b}=[0.97,0.89,0.85,0.37,0.35,0.31,0.25,0.22,0.18]$ and $\text{scores}_{a}=[0.80,0.72,0.65,0.46,0.39,0.28]$. Suppose the query pair $p$ has a score $s(p)=0.34$ and belongs to the b group. The algorithm sets $g=b$ and $\text{pos}=0$. It then performs a linear search to find the position of $0.34$ within $\text{scores}_{b}$, which falls between $0.31$ and $0.35$, so $\text{pos}_{b}=6$. Next, $\alpha$ is calculated as the proportion of minority scores, $\alpha=|\text{scores}_{a}|/|D|=6/15=0.4$, and the quantile $q=\text{pos}_{g}/|\text{scores}_{g}|=6/9\approx 0.67$. Using $q$, we compute the corresponding position in the minority scores: $\text{pos}_{a}=\left\lceil q\times|\text{scores}_{\bar{g}}|\right\rceil=0.67\times 6\approx 4$, giving the score $0.46$ in the minority group. Finally, we compute a weighted average of the scores $0.31$ and $0.46$ using weights $\alpha$ and $1-\alpha$, resulting in the calibrated score $\hat{s}(p)$ = $0.4\times 0.46$ + $0.6\times 0.31$ = $0.37$ Thus, the algorithm returns a calibrated score of $0.37$ for the query pair $p$. The steps are illustrated in Figure 2. $\blacksquare$
IV-B Theoretical guarantees
Algorithm 1 returns a calibrated score for any given query pair, and therefore, in essence, it computes a score function $\hat{s}$. In the following theorem, $s$ is the input score function, $\hat{s}$ is the calibrated score function, and $s^{*}$ is the optimal matching score function (Definition 2). Let $n_{a}$ and $n_{b}$ denote the number of the minority and majority samples in $D$. The theorem provides an upper bound on the bias of $\hat{s}$ and asymptotically upper bounds its risk, limiting $\hat{s}$ deviation from $s^{*}$.
@@ IV-B Theoretical guarantees
For the calibrated score $\hat{s}$ obtained from Algorithm 1, with input score $s$ and optimal score function $s^{*}$, the following asymptotic bounds hold, where $n=\min(n_{a},n_{b})$:
$\textit{bias}(\hat{s},\text{PR})=O(n^{-1}),$ (6)
$\textit{risk}(s^{*},\hat{s})=O(\log(n)^{-1/2}).$ (7)
@@
To prove Eq. 6, we start from the left-hand side:
[…]
Here, $P_{g}(\hat{S})=P(\hat{s}(X)=\hat{S}|A=g),g\in\{a,b\}$ and $P$ is the barycenter of $P^{s}_{a}$ and $P^{s}_{b}$. The integral of a function over $[0,1]$ is always bounded by the function’s supremum on $[0,1]$. Thus, from Eq. 8, we deduce:
$\textit{bias}(\hat{s},\text{PR})\leq\sup_{\theta\in[0,1]}\big{|}P_{a}(\hat{S}\leq\theta)-P_{b}(\hat{S}\leq\theta)\big{|}$ (9)
Since $P$ is a barycenter obtained from the quantile-based method; we write the right-hand side of Eq. 9 as follows:
$\textit{bias}(\hat{s},\text{PR})\leq\sup_{\theta\in[0,1]}\big{|}P(\mathbb{F}_{a}(s(X))\leq\theta)-P(\mathbb{F}_{b}(s(X))\leq\theta)\big{|}$ (10)
where $\mathbb{F}_{g}$ is the CDF of the scores for group $g$. This means an estimation of $P(\mathbb{F}_{g}(s(X))\leq\theta)$ is $m_{g}/(n_{g}+1)$ where $m_{g},g\in\{a,b\}$ is the number of pairs $p$ in $D$ that are from group $g$ and $s(p)<\theta$. Therefore, we write Eq. 10 as
$\textit{bias}(\hat{s},\text{PR})\leq\sup_{\theta\in[0,1]}\big{|}\frac{m_{a}}{n_{a}+1}-\frac{m_{b}}{n_{b}+1}\big{|}$ (11)
Since $0\leq m_{g}\leq n_{g}$, the difference is maximized when one group has $m_{g}=0$ and the other $m_{\bar{g}}=n_{g}$, yielding a worst‐case difference rate of $O(1/\min(n_{a},n_{b}))=O(n^{-1})$.
The proof of Eq. 7 relies on the fact that the score $\tilde{s}$, obtained from the average barycenter of $s_{a}$ and $s_{b}$, has the minimal risk. The barycenter minimizes risk by reducing the total Wasserstein distance. Since risk is defined as the expected difference between the calibrated score and the original score, it is directly tied to the transport cost in the Wasserstein distance. The barycenter ensures that adjustments to align group-specific score distributions are as small as possible. The cost function $|x-y|^{2}$ in the Wasserstein distance quantifies the effort required to modify scores, which directly affects the expected deviation. By minimizing this cumulative transport cost, the barycenter produces a calibrated score that reduces bias while making minimal changes to the original scores. This property makes it a theoretically optimal solution for balancing fairness and accuracy in score calibration.
The proof uses the following concentration inequality, commonly applied in statistical learning theory [32, 33]:
$P\left(|\tilde{s}(X)-\hat{s}(X)|\geq\delta\right)\leq c\exp\left(-Cb_{n}\delta^{2}\right),$ (12)
where $c$ and $C$ are constants, $b_{n}$ depends on the sample size, and $\exp(-Cb_{n}\delta^{2})$ reflects the rapid decay of probability as $\delta$ grows. Integrating this inequality over all possible deviations bounds the risk:
[…]
Substituting the tail probability and solving the integral gives:
$E\left[|\tilde{s}(X)-\hat{s}(X)|\right]\leq\frac{c\sqrt{\pi}}{2\sqrt{Cb_{n}}}=O(b_{n}^{-1/2}).$ (14)
For deep learning models used in classification, including those applied to matching in this work, the rate parameter $b_{n}$ is typically $\log(n)$ due to the high capacity of neural networks and implicit regularization effects [34, 35, 36]. Substituting $b_{n}=\log(n)$ yields the upper bound $O(\log(n)^{-1/2})$. ∎
Theorem 1 shows that as the sample size increases, the bias of $\hat{s}$ vanishes and its risk reaches the minimum value. As a result, the calibrated score function $\hat{s}$ converges to the optimal score function $s^{*}$, achieving both fairness and accuracy. Also, this theorem specifies what “sufficiently large” dataset $D$ means in Section IV-A. From Eq. 6, if the maximum allowable bias after calibration is $\epsilon$, then the dataset must satisfy $|D|\geq O\left(\frac{1}{\epsilon}\right)$.
V Conditional Calibration
@@ 1 $\text{scores}_{b}\leftarrow[]$; $\text{scores}_{a}\leftarrow[]$;
2 /* Find optimal decision threshold */
3 $\gamma\leftarrow\text{meanshift}(D)$;
@@ 12             $\text{size}\leftarrow\text{size}+1$;
14
15sort($\text{sort}(\text{scores}_{a})$; $\text{sort}(\text{scores}_{b})$; $g\leftarrow p[A]$; $\text{pos}_{g}\leftarrow 0$;
@@ 17       $\text{pos}_{g}\leftarrow\text{pos}_{g}+1$;
[…]
The core idea behind Algorithm 2, is to include predicted labels using variables $matched$ and $matched_{i}$. Here, $matched$ is the predicted label for the query pair $p$, while $matched_{i}$ is the predicted label for each pair in $D$. This is illustrated in Lines 2 and 2 of Algorithm 2. To estimate labels- whether for the query pair or pairs in $D$- we require a decision threshold to assign predicted labels based on scores. In traditional binary classification, this threshold is determined on a labeled validation set to optimize a performance metric (e.g., F1-score) and then applied to unlabeled test data. In our case, we only have a query pair and a set $D$ of observed pairs. To choose an effective decision threshold, we use a non-parametric clustering algorithm, meanshift [37], as shown in Line 2. The meanshift algorithm places a window around each data point, computes the mean within the window, and shifts the window toward the mean until convergence. This process clusters points around the modes of the distribution. When applied to score values, which tend to cluster near 0 and 1 (which is the case for record matching problem), the algorithm identifies two clusters. The decision threshold, denoted as $\gamma$, is then set as the midpoint between the centers of these clusters, effectively separating the data into two classes without the need for labeled data.
The parameter $\gamma$ has two purposes in C-Calib. In Line 2, it estimates the label of query pair, and in Line 2, it separates the samples in $D$ into positive and negative samples, assuming $D$ is unlabeled. Unlike decision threshold of $\theta$ in the final down-stream applications, $\gamma$ is specific to the calibration algorithm and does not decide the final label of $p$. While $\gamma$ might appear inconsistent with a threshold-agnostic algorithm, it acts as a tuning parameter rather than a classification threshold. By “threshold-agnostic,” we mean the algorithm mitigates bias across all classification thresholds in the final results.
Moving forward, using the conditional check in Line 2, the algorithm adjusts scores based on the query pair’s predicted label. If the query pair is predicted as positive, only pairs in $D$ predicted as positive are included in the calibration; the same applies for a negative prediction. The remainder of the algorithm follows the process of Algorithm 1 described in Section IV-A. We call this approach “conditional” because it calibrates only among pairs sharing the same predicted label as the query pair, consequently enforcing EOD and EO. Now, we demonstrate the entire process with the following example.

### v3

@@ 5 $\text{append}(\text{scores}_{g},s(p_{i})+\mathcal{N}(0,\sigma^{2}))$;
[…]
@@ 11 $\text{pos}_{g}\leftarrow\text{pos}_{g}+1$;
14 if $g=a$ then
15 $\text{pos}_{a}\leftarrow\text{pos}_{g}$;
16 $\text{pos}_{b}\leftarrow\left\lceil q\times|\text{scores}_{b}|\right\rceil$; /* Cross-group map */
17
18else
19 $\text{pos}_{b}\leftarrow\text{pos}_{g}$;
20 $\text{pos}_{a}\leftarrow\left\lceil q\times|\text{scores}_{a}|\right\rceil$; /* Cross-group map */
21
@@ 19 $\text{pos}_{b}\leftarrow\text{pos}_{g}$;
The second point concerns the input dataset $D=\{p_{1},\dots,p_{|D|}\}$, used to estimate $P_{a}$ and $P_{b}$. This dataset does not require ground-truth labels–only matching scores are needed, which can be obtained by applying the score function $s$. As a result, $s$ can be calibrated as a black box without access to the training data used to learn $s$, such as in deep learning-based matchers. This significantly broadens the applicability of the calibration method. The only assumption is that the pairs in $D$ are i.i.d. samples from the space of possible record pairs and that the dataset is large enough to reliably estimate the score distributions. The primary goal of the algorithm is to calibrate the scores for one or more query pairs. When the number of query pairs is large enough (as discussed in Section 5.2), the dataset $D$ can simply be the set of query points itself. We examine the effect of choosing a dataset $D$ with a distribution different from that of the query points in Section 6.2.5. We now describe each step of Algorithm 1.
@@
To calibrate the score for a query pair $p$, the algorithm identifies the group $g$ of $p$ (Line 1) and initializes a pospos variable within the sorted score list (Line 1). It then iterates through the sorted scores to locate the position where $s(p)$ would fit, counting scores greater than $s(p)$ within group $g$ (Lines 11). This position, relative to the list length, defines a quantile $q$, representing the percentage of scores in group $g$ greater than $s(p)$ (Line 1). The algorithm then calculates $\text{pos}_{\bar{g}}=\left\lceil q\times|\text{scores}_{\bar{g}}|\right\rceil$, where $\bar{g}$ is the opposite group (i.e., if $g=a$, then $\bar{g}=b$, and vice versa). This position is used to find the score in the other group’s list that corresponds to the same quantile $q$ as the query score in its own group.
@@
Consider dataset $D$ shown in Figure 2, containing 15 record pairs scored by $s$ and labeled a or b by a sensitive attribute. The pairs and scores are: $(0.45,a)$, $(0.82,a)$, $(0.90,b)$, $(0.71,a)$, $(0.84,b)$, $(0.67,a)$, $(0.38,b)$, $(0.98,b)$, $(0.36,b)$, $(0.38,a)$, $(0.32,b)$, $(0.29,a)$, $(0.24,b)$, $(0.21,b)$, $(0.19,b)$. Following the calibration algorithm, Gaussian noise $\mathcal{N}(0,0.05^{2})$ is added to each score, resulting in the following (rounded) values: $(0.46,a)$, $(0.80,a)$, $(0.89,b)$, $(0.72,a)$, $(0.85,b)$, $(0.65,a)$, $(0.37,b)$, $(0.97,b)$, $(0.35,b)$, $(0.39,a)$, $(0.31,b)$, $(0.28,a)$, $(0.25,b)$, $(0.22,b)$, $(0.18,b)$. Sorting these in descending order gives $\text{scores}_{b}=[0.97,0.89,0.85,0.37,0.35,0.31,0.25,0.22,0.18]$ and $\text{scores}_{a}=[0.80,0.72,0.65,0.46,0.39,0.28]$. Suppose the query pair $p$ has a score $s(p)=0.34$ and belongs to the b group. The algorithm sets $g=b$ and $\text{pos}=0$. It then performs a linear search to find the position of $0.34$ within $\text{scores}_{b}$, which falls between $0.31$ and $0.35$, so $\text{pos}_{b}=6$. Next, $\alpha$ is calculated as the proportion of minority scores, $\alpha=|\text{scores}_{a}|/|D|=6/15=0.4$, and the quantile $q=\text{pos}_{g}/|\text{scores}_{g}|=6/9\approx 0.67$. Using $q$, we compute the corresponding position in the minority scores: $\text{pos}_{a}=\left\lceil q\times|\text{scores}_{\bar{g}}|\right\rceil=0.67\times 6\approx 4$, giving the score $0.46$ in the minority group. Finally, we compute a weighted average of the scores $0.31$ and $0.46$ using weights $\alpha$ and $1-\alpha$, resulting in the calibrated score $\hat{s}(p)$ = $0.4\times 0.46$ + $0.6\times 0.31$ = $0.37.$ Thus, the algorithm returns a calibrated score of $0.37$ for the query pair $p$. The steps are illustrated in Figure 2. $\blacksquare$
4.2 Theoretical Analysis of Calib
Our analysis follows standard arguments from the optimal-transport and fair-regression literature for 1D distributional alignment and Wasserstein barycenters [39, 31, 10]. Algorithm 1 returns, for each input pair, a calibrated score and therefore defines a new score function $\hat{s}$. In this section, $s$ denotes the original score function, $\hat{s}$ the calibrated score, and $s^{*}$ the optimal fair score in the sense of Definition 2. Let $n_{a}$ and $n_{b}$ be the numbers of calibration samples in each group, and let $n=\min(n_{a},n_{b})$.
We state mild assumptions used in our analysis in this section. Intuitively, these assumptions ensure that the score distributions behave “smoothly,” so that quantiles are uniquely defined and the empirical version of Calib is a stable approximation of its population counterpart. Technically, they allow us to apply standard empirical-process tools (e.g., Dvoretzky–Kiefer–Wolfowitz and quantile-process convergence) and guarantee that the Wasserstein barycenter map is well-defined and unique in one dimension. Importantly, these conditions are not restrictive in practice: modern matching or similarity models produce real-valued scores that can always be rescaled to $[0,1]$, and introducing an arbitrarily small jitter has no effect on application-level decisions—its role is purely to avoid pathological ties in the theoretical analysis.
The following are the assumptions.
  • Assumption 1: For each group $g\in\{a,b\}$, the random variable $S_{g}$ (i.e., $s(X)$ for $A=g$) has an absolutely continuous distribution on $[0,1]$ with CDF $F_{g}$ and quantile function $Q_{g}=F_{g}^{-1}$. The two distributions share the same support $[0,1]$, ensuring that both groups can be transported to a common barycenter in a well-defined way.
  • Assumption 2: The calibration samples are i.i.d., and Algorithm 1 adds an independent Gaussian jitter to break ties, ensuring that empirical CDFs and quantiles are well-defined and converge uniformly to their population counterparts.
Under these assumptions, the population version of Calib maps each group to the Wasserstein–2 barycenter of the two group-wise score distributions. Let $\alpha=\Pr(A=a)$ and define the barycenter quantile
$\tilde{Q}(p)=\alpha\,Q_{a}(p)+(1-\alpha)\,Q_{b}(p),\qquad p\in[0,1].$
The corresponding population barycenter mapping is
$\tilde{s}(x)=\tilde{Q}(F_{A(x)}(s(x))),$
which ensures that $\tilde{s}(X)\mid A=a$ and $\tilde{s}(X)\mid A=b$ have identical distributions. As standard in 1D optimal transport, $\tilde{s}$ uniquely minimizes the risk among all DP-fair score mappings and therefore coincides with $s^{*}$.
Algorithm 1 computes the empirical analogue of this map by replacing each $F_{g}$ and $Q_{g}$ with their empirical counterparts. The following theorem shows that this empirical map converges to $s^{*}$ at the parametric rate.
@@ 4.2 Theoretical Analysis of Calib
Let $\hat{s}$ be the calibrated score produced by Algorithm 1 and let $s^{*}$ denote the population barycenter score defined above. Under Assumptions 1–2,
$\textit{bias}(\hat{s},\text{PR})=O_{p}(n^{-1/2}),\qquad\textit{risk}(s^{*},\hat{s})=O_{p}(n^{-1/2}),$
where $n=\min(n_{a},n_{b})$.
@@
Bias bound: For any score function $s$, recall that
$\text{PR}_{g}(s,\theta)=\Pr(s(X)\geq\theta\mid A=g)=1-F_{g}^{s}(\theta),$
where $F_{g}^{s}$ is the CDF of $s(X)\mid A=g$. Hence
$\big|\text{PR}_{a}(\hat{s},\theta)-\text{PR}_{b}(\hat{s},\theta)\big|=\big|F_{a}^{\hat{s}}(\theta)-F_{b}^{\hat{s}}(\theta)\big|.$
Thus
$\textit{bias}(\hat{s},\text{PR})=\int_{0}^{1}\big|F_{a}^{\hat{s}}(\theta)-F_{b}^{\hat{s}}(\theta)\big|\,d\theta.$
In the population, both groups are mapped to the same barycenter distribution, so $F_{a}^{\tilde{s}}=F_{b}^{\tilde{s}}$. With finite samples, Algorithm 1 replaces $F_{g}$ by the empirical CDF $\hat{F}_{g}$. By the Dvoretzky–Kiefer–Wolfowitz inequality,
$\sup_{\theta\in[0,1]}|\hat{F}_{g}(\theta)-F_{g}(\theta)|=O_{p}(n_{g}^{-1/2}),\qquad g\in\{a,b\}.$
Since the population barycenter CDFs coincide, the sup-norm difference between the two empirical calibrated CDFs is bounded by the sum of their uniform deviations:
$\sup_{\theta\in[0,1]}\big|F_{a}^{\hat{s}}(\theta)-F_{b}^{\hat{s}}(\theta)\big|=O_{p}(n^{-1/2}).$
Integrating over $\theta$ preserves this rate, giving
$\textit{bias}(\hat{s},\text{PR})=O_{p}(n^{-1/2}).$
Risk bound: The risk satisfies
$\textit{risk}(s^{*},\hat{s})={\tt\mathbb{E}}\!\left[\,|s^{*}(X)-\hat{s}(X)|\,\right].$
Since $s^{*}=\tilde{s}$, it suffices to control the deviation between the empirical barycenter and the population barycenter in quantile space.
Let $Q_{g,n}$ be the empirical quantile of group $g$. The empirical barycenter quantile is
$\tilde{Q}_{n}(p)=\alpha\,Q_{a,n}(p)+(1-\alpha)\,Q_{b,n}(p).$
Standard quantile-process theory gives
${\tt\mathbb{E}}\!\left[\,|Q_{g,n}(U)-Q_{g}(U)|\,\right]=O(n_{g}^{-1/2}),\qquad g\in\{a,b\},$
where $U\sim\mathrm{Unif}[0,1]$. Hence
${\tt\mathbb{E}}\!\left[\,|\tilde{Q}_{n}(U)-\tilde{Q}(U)|\,\right]=O_{p}(n^{-1/2}).$
Mapping back through $F_{A(x)}$ preserves the order, so
$\textit{risk}(s^{*},\hat{s})=O_{p}(n^{-1/2}).$
Thus both DP bias and risk converge at the parametric rate $O_{p}(n^{-1/2})$. We use $O_{p}(\cdot)$ because the bounds involve random empirical CDFs and quantiles whose fluctuations shrink at rate $n^{-1/2}$ only in probability. ∎
Theorem 1 shows that as $n$ grows, the calibrated score $\hat{s}$ approaches the optimal DP-fair score $s^{*}$, simultaneously achieving fairness and minimizing deviation from the original scores.
5 Conditional Score Calibration
@@ 1 $\text{scores}_{b}\leftarrow[]$; $\text{scores}_{a}\leftarrow[]$;
2 $\gamma\leftarrow\text{meanshift}(D)$; /* Estimate decision threshold */
@@ 11    $\text{size}\leftarrow\text{size}+1$;
[…]
@@ 16 $\text{pos}_{g}\leftarrow\text{pos}_{g}+1$;
Algorithm 2 $\texttt{C-Calib}(D,s,p)$
5.1 C-Calib Algorithm
The core idea behind Algorithm 2 is to incorporate predicted labels using variables match and $\text{match}_{i}$. Here, match is the predicted label for the query pair $p$, while $\text{match}_{i}$ is the predicted label for each pair in $D$. This is illustrated in Lines 2 and 2 of Algorithm 2. To estimate labels—whether for the query pair or pairs in DD—we require a decision threshold to assign predicted labels based on scores. In traditional binary classification, this threshold is determined on a labeled validation set to optimize a performance metric (e.g., F1-score) and then applied to unlabeled test data. In our case, we only have a query pair and a set $D$ of observed pairs. To choose an effective decision threshold, we use a non-parametric clustering algorithm, meanshift [11], as shown in Line 2. The meanshift algorithm places a window around each data point, computes the mean within the window, and shifts the window toward the mean until convergence. This process clusters points around the modes of the distribution. When applied to score values, which tend to cluster near 0 and 1 (which is the case for record matching problem), the algorithm identifies two clusters. The decision threshold, denoted as $\gamma$, is then set as the midpoint between the centers of these clusters, effectively separating the data into two classes without the need for labeled data.
The parameter $\gamma$ has two purposes in C-Calib. In Line 2, it estimates the label of query pair, and in Line 2, it separates the samples in $D$ into positive and negative samples, assuming $D$ is unlabeled. Unlike decision threshold of $\theta$ in the final down-stream applications, $\gamma$ is specific to the calibration algorithm and does not decide the final label of $p$. While $\gamma$ might appear inconsistent with a threshold-independent algorithm, it acts as a tuning parameter rather than a classification threshold. By “threshold-independent,” we mean the algorithm mitigates bias across all classification thresholds in the final results.
Moving forward, using the conditional check in Line 2, the algorithm adjusts scores based on the query pair’s predicted label. If the query pair is predicted as positive, only pairs in $D$ predicted as positive are included in the calibration; the same applies for a negative prediction. The remainder of the algorithm follows the process of Algorithm 1 described in Section 4.1. We call this approach “conditional” because it calibrates only among pairs sharing the same predicted label as the query pair, consequently enforcing EOD and EO. Now, we demonstrate the entire process with the following example.